Two airlines compete in prices. The can differentiate the product so that their respective demand functions look like

Qa=4000-25Pa+12Pb

Qb=3000-20PPa

AVCa=MCa=160

AVCb=MCb=180

profit functions are

πa(Pa,Pb)=(Pa-160)(4000-25Pa+12Pb)

πb(Pa,Pb)=(Pb-180)(3000-20Pb-10Pa

Reaction functions are

Pa=160+(6/25)Pb

Pb=165+(1/4)Pa

(2 points) Use this information to draw each firms reaction function. (Hint: let pA be on the horizontal axis and pB in the vertical axis)
(3 points) Find the Nash equilibrium choice of prices.
Answer #1

a)

Pb Pa=160+(6/25)*Pb
0.00 160.00
50.00 172.00
100.00 184.00
150.00 196.00
200.00 208.00
218.09 212.34
250.00 220.00
Pb=165+(1/4)*Pa Pa
165.00 0.00
177.50 50.00
190.00 100.00
202.50 150.00
215.00 200.00
218.09 212.34
227.50 250.00

b)

We are given

Pb=165+(1/4)*Pa

Set Pa=160+(6/25)*Pb

Pb=165+(1/4)*[160+(6/25)*Pb]=165+40+(3/50)Pb

(47/50)Pb=205

Pb=(205*50/47)=$218.0851 or say $218.09

Pa=160+(6/25)*Pb=160+(6/25)*218.0581=$212.3409 or say $212.34